OpenStudy (anonymous):

Multiply or divide as indicated. x^10/x^4

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use this : \(\large \frac{x^m}{x^n} = x^{m-n}\)

ganeshie8 (ganeshie8):

\(\large \frac{x^{10}}{x^4} = x^{10-4} = x^6\)